OpenStudy (anonymous):

The area of a large circle is 9 times the area of a small circle. How many times as long is the radius of the large circle compared with the radius of the small circle?

OpenStudy (accessdenied):

I would use the ratio of areas and take the positive square root to find the ratio of dimensions. Ratio of Areas: 9A/A = 9/1

OpenStudy (anonymous):

3 times! to find the area of a circle the radius is r^2 so: pi*r^2=9*pi*R^2 if you solve it you'll find r=3R

OpenStudy (anonymous):

Alright, thank you both :)

OpenStudy (accessdenied):

You're welcome. :)